OpenStudy (anonymous):

Solve the equation: –5 = s – 3

OpenStudy (saifoo.khan):

s = -5 + 3 s = -2

OpenStudy (anonymous):

3-s=-5 -5+3 s=2

OpenStudy (anonymous):

s=-2